# X-Fi: A Modality-Invariant Foundation Model for Multimodal Human Sensing

### v2

@@ A.2 Evaluation of modality-specific features combining methods
Our objective is to assess whether concatenation is the most effective method for combining modality-specific features into multi-modal embedding. To investigate this, we conducted a controlled experiment on the MM-Fi HPE task. In this experiment, we compared the concatenation method with the feature addition approach, where the feature values from each modality are summed and then divided by the number of modalities. As demonstrated in the Table 7, the feature addition approach failed to preserve distinctive modality features in the multimodal embeddings, resulting in average performance deteriorations of 15.7 mm on MPJPE and 5.9 mm on PA-MPJPE.
@@
A.3 Ablation Studies
To demonstrate the effectiveness of our proposed modality-invariant foundation model and the corresponding optimization process, we evaluate the impact of dropout layer, positional encoding,and the optimizer choice. And the results are illustrated in Table 67.
To evaluate the effect of dropout layers, we incorporated a dropout layer with a rate of 0.3 into all transformer structures within X-Fi and assessed its performance on the MMFi dataset for the HAR task. The experimental results indicated that the introduction of the dropout layer led to a 4.6% decrease in averaged accuracy. Furthermore, the dropout layer significantly impaired the model’s predictive performance, particularly when handling challenging modalities.
To assess the impact of different optimizer choices on experimental results, we conducted a set of controlled experiments using the SGD optimizer with a learning rate scheduler on the MM-Fi HPE task. The results indicated that the AdamW optimizer facilitated better convergence to a more optimal local minimum, leading to an average performance improvement of xx mm on MPJPE and x mm on PA-MPJPE.

### v3

@@ A.2 Evaluation of modality-specific features combining methods
Our objective is to assess whether concatenation is the most effective method for combining modality-specific features into multimodal embedding. To investigate this, we conducted a controlled experiment on the MM-Fi HPE task. In this experiment, we compared the concatenation method with the feature addition approach, where the feature values from each modality are summed and then divided by the number of modalities. As demonstrated in the Table 11, the feature addition approach failed to preserve distinctive modality features in the multimodal embeddings, resulting in average performance deteriorations of 15.7 mm on MPJPE and 5.9 mm on PA-MPJPE.
@@
A.3 Comparison with SOTA multimodal fusion approaches
To demonstrate the the multimodal fusion effectiveness of our proposed X-Fi, we conduct comparisons with SOTA multimodal fusion approaches that are based on fixed modality combinations, e.g. ImmFusion (Chen et al., 2023) and Meta-Transformer(Zhang et al., 2023). The experiments are conducted on the MM-Fi dataset(Yang et al., 2024) for the HPE task. For ImmFusion, we first train the model using the modality combination from the original paper (I+D+R) and then adapt the method to include all modalities (I+D+R+L+W). For Meta-Transformer, we use the Meta-Transformer-B16 encoder as the backbone. Initially, we fine-tune the pretrained encoder weights (from the original repository) and train a regression head on all modalities (I+D+R+L+W). Next, we freeze the encoder weights and fine-tune the regression head on Rgb+Depth+LiDAR modalities, obtaining the Meta-Transformer-B16F(I+D+L) result. Finally, we fine-tune both the encoder and the regression head on the three modalities to produce the Meta-Transformer-B16T(I+D+L) result. The results shown in the Table 7 demonstrate X-Fi’s superior performance.
Modalities Methods MPJPE (mm) PA-MPJPE (mm)
Modalities Methods MPJPE (mm) PA-MPJPE (mm)
I + D + L Meta-Transformer-B16F 105.2 70.3
Meta-Transformer-B16T 96.2 68.1
X-Fusion 84.8 48.2
I + D+ R ImmFusion 207.3 146.8
X-Fusion 83.4 47.3
I + D + L + R + W ImmFusion 107.9 63.7
Meta-Transformer-B16T 91.4 64.9
X-Fusion 83.7 47.6
Table 7: Comparison between X-Fi and SOTA multimodal fusion approaches, e.g. ImmFusion and Meta-Transformer.
A.4 Ablation Studies
To demonstrate the effectiveness of our proposed modality-invariant foundation model and the corresponding optimization process, we evaluate the impact of the dropout layer, positional encoding, Key-Value generators, the optimizer choice, the number of iterations on X-Fusion block, and the number of layers in cross-modal transformer. The results are illustrated in the Table 68,& 11.
Dropout Layer.
To evaluate the effect of dropout layers, we incorporated a dropout layer with a rate of 0.3 into all transformer structures within X-Fi and assessed its performance on the MMFi dataset for the HAR task. The experimental results indicated that the introduction of the dropout layer led to a 4.6% decrease in average accuracy. Furthermore, the dropout layer significantly impaired the model’s predictive performance, particularly when handling challenging modalities.
Optimizer Choice.
To assess the impact of different optimizer choices on experimental results, we conducted a set of controlled experiments using the SGD optimizer with a learning rate scheduler on the MM-Fi HPE task. The results indicated that the AdamW optimizer facilitated better convergence to a more optimal local minimum, leading to an average performance improvement of 50.9 mm on MPJPE and 44.6 mm on PA-MPJPE.
Positional Encoding.
@@ A.4 Ablation Studies
Key-Value Generators.
To assess the impact of Key-Value Generators, we conducted an experiment by removing the Key-Value Generators and directly use modality-specific features to inject information in the cross-attention module. The ablation study results shown in the Table 8 proves that utilizing key-value generators to obtain key-value pairs is more efficient than directly using initial modality features as k-v.
Ablation Studies
on XRF55 HAR
Number of Iterations
Key-Value Generators
(with/without)
Depth of cross-modal
transformer (#)
Methods # = 2 # = 4 without with # = 4 # = 1
R 80.6% 83.9% 82.0% 83.9% 82.8% 83.9%
W 24.2% 55.7% 26.8% 55.7% 54.8% 55.7%
RF 38.8% 42.5% 40.0% 42.5% 40.5% 42.5%
R + W 84.1% 88.2% 85.0% 88.2% 87.1% 88.2%
R + RF 84.1% 86.5% 84.7% 86.5% 84.2% 86.5%
W + RF 48.0% 58.1% 50.2% 58.1% 54.6% 58.1%
R + W + RF 86.3% 89.8% 86.7% 89.8% 87.2% 89.8%
Table 8: Ablation studies on Key-Value generators, number of iterations on X-Fusion block, and number of layers in cross-modal transformer.
Number of iterations on X-Fusion block.
To assess the impact of the number of iterations on X-Fusion block, we conducted an experiment by adjusting the number. The ablation study results shown in the Table 8 demonstrate that when the number of iterations is set to 2, the model finds difficulties in digging and preserving features from weak modalities. We need to set the number of iterations on the X-Fusion block to at least 4 to repeatedly inject information from weaker modalities, enabling the model to more robustly fit all participating modalities during training.
Number of layers in cross-modal transformer.
To assess the impact of the number of layers in cross-modal transformer, we have added the ablation of the number of the layers (depth) in cross-modal transformers on the XRF55 HAR task. The default depth in our experiments is 1 and the depth we choose in the comparison experiment is 4. The results shown in the Table 8 indicate that adding the number of the layers in the cross-modal transformer will not yield better performance.
A.5 HPE quantitative result investigation
In order to have a more detailed look at the HPE quantitative result and show the observation that RGB modality captures better leg movements and depth modality is better for the arm, we have zoomed in the areas and the picture after enlargement has been added in the Figure 5. To seek the reason for such observation, we explore the characteristics of the rgb data and the depth data in MM-Fi. We find that due to the arm’s movement typically being faster than the legs, combined with the insufficient frame rate of the camera, the arms in the RGB samples are often blurry, while the legs appear very clear. This results in poor RGB-based estimation performance for the arms. In contrast, in depth images, the arm contours are very clear, but the legs almost blend into the background, making their contours indistinct, which leads to poor depth-based estimation performance for the legs. We also include two images samples from rgb and depth in the Figure 6.
Figure 5: The detailed comparison among the rgb HPE result, the depth HPE result, and the ground truth.
Figure 6: The comparison between a rgb training sample and a depthtraining sample.
A.6 Continual Learning on X-Fi
In order to validate the continual learning performance of X-Fi, we carry out experiment which gradually add new modality each time. We first trained an X-Fi model from scratch with RGB and Depth modalities. Then we added mmWave modality and fine-tuned the model based on the pretrained parameters. The results is shown in the Table 9.
MM-Fi MPJPE (mm) PA-MPJPE (mm)
Methods
train with Rgb
and Depth
Fine-tune
with mmWave
train with Rgb
and Depth
Fine-tune
with mmWave
I 117.9 105.2 71.7 64.3
D 108.5 130.2 56.2 52.8
R - 128.1 - 72.4
I + D 99.2 106.4 59.2 53.6
I + R - 92.8 - 58.7
D + R - 115.7 - 50.3
I + D + R - 100.8 - 51.8
Table 9: Continual Learning performance of X-Fi.
Model Structure Number of Parameters
X-Fusion block
(5 modalities)
17,903,667
X-Fusion block
(4 modalities)
14,748,723
X-Fusion block
(3 modalities)
11,593,779
X-Fusion block
(2 modalities)
8,438,835
X-Fusion block
(1 modality)
5,283,891
VO Transformer
fusion block
12,636,723
Table 10: Comparison of number of parameters used in VO Transformer fusion block and X-Fusion block.
@@ Methods
Figure 7: The detailed architecture of the proposed key-value generators and cross-attention modules.
Figure 8: The illustration of the training processes we experimented with: cross-attention-based dimension reduction (left image) and the addition of contrastive loss (right image).